activity of 72 200Au of mass 3x10-6 g is 58.9 ci.its half life period is ........ 


The ans is 28.8x100 s.How.

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

number of atoms=(3*10-6)/200 =a,activity=-da/dt=λa, (lambda is decay constant)=58.9*3.7*10^10=b,λ=b/a,half life=ln2/λ